Carlo is half as old as Raul. James is 6 years older than Carlo. Four years ago, the sum of their ages was 50. Find their present age.

Answer by Boreal(15235)

C=x
R=2x
J=x+6
4 years ago, they were
x-4, 2x-4,x+2
That sum, 4x-6, is 50
4x-6=50
4x=46
x=14
Carlo 14
Raul is 28
James is 20
Now the age sum is 62
4 years ago it was 62-(3*4)=50